OpenStudy (anonymous):

The rate at which water is pumped out of a dam, in L/min, t minutes after the pump is started is (shown in next post). a) How much water is pumped out in the 40th minute? b) Find the volume of water pumped out at any time, t, after the pump is started. c) How much water is pumped out after 40 minutes? d) Find the average rate at which water is pumped in the first hour e) How long would it take to fill a tank holding 1600 litres? Can you do your explanation using drawings and Equations please?

OpenStudy (anonymous):

\[\frac{ dV }{ dt } = 5 + \cos \frac{ \pi t }{ 40 }\]

OpenStudy (anonymous):

@jim_thompson5910 @Luigi0210